OpenStudy (anonymous):

use the quadratic formula to solve the equation -x^2+7x=5

OpenStudy (anonymous):

A. \[-\frac{ 7 }{ 2 }\pm \frac{ \sqrt{29} }{ 2 }\] B. \[\frac{ 7 }{ 2 }\pm \frac{ \sqrt{29} }{ 4}\] C. \[\frac{ 7 }{ 2 }\pm \frac{ \sqrt{29} }{ 2 }\] D. \[\frac{ 7 }{ 4 }\pm \frac{ \sqrt{29}}{ 2 }\]

OpenStudy (mathmale):

Hi, Rockie, How about subtracting 5 from both sides of the equation? Can you then identify the values of the coefficients a, b and c of the resulting quadratic equation in standard form?

OpenStudy (anonymous):

so it would be -x^2+7x-5 then what else do you do ?

OpenStudy (mathmale):

Rockie, please identify the values of the three coefficients: The coefficient of x^2 is -1, so a = -1 That of x is 7, so b = 7 The constant term is -5, so c = -5. Remember the quadratic formula? -b plus or minus Sqrt(b^2 - 4ac) x=---------------------------- 2a Substitute the values of your a, b and c into this formula. Compare your results to the four possible answers which you've shared.

OpenStudy (anonymous):

im so confused o.O

OpenStudy (mathmale):

Rockie, I'm sure you've seen the quadratic formula before. It's almost always written in terms of 3 variables, a, b and c. You have values for a, b and c, so please substitute them into the quadratic formula to obtain the x values which are the solutions you've been seeking.

OpenStudy (anonymous):

okay i did it and i got B is that right ?